OpenStudy (anonymous):

Write the equation of the circle centered at (0, 0) that has a radius of r=6. Top Half: y= Bottom Half: y=

OpenStudy (anonymous):

x^2+y^2=6^2

OpenStudy (anonymous):

looks like you want to solve for y

OpenStudy (anonymous):

y^2=36-x^2 top half y=sqrt(36-x^2) bottom half y=-sqrt(36-x^2)

OpenStudy (anonymous):

Awesome thanks! I had that equation down but I didn't know what was next. You rock!

OpenStudy (anonymous):

yw